# Eigenvalue bounds for Schrödinger operators with complex potentials on compact manifolds

### v1

@@ Abstract.
We prove eigenvalue bounds for Schrödinger operator $-\Delta_{g}+V$ on compact manifolds with complex potentials $V$. The bounds depend only on an $L^{q}$-norm of the potential, and they are shown to be optimal, in a certain sense, on the round sphere and more general Zoll manifolds. These bounds are natural analogues of Frank’s [9] results in the Euclidean case.
@@ 1.1. Spectral inclusion
with a complex-valued potential $V$. The unperturbed operator $-\Delta_{g}$ is a non-negative, unbounded self-adjoint operator in $L^{2}(M)$, defined with respect to the Riemannian volume measure ${\rm d}x$. Since $-\Delta_{g}$ has compact resolvent, its spectrum consists of nonnegative eigenvalues $0=\lambda_{0}^{2}\leq\lambda_{1}^{2}\leq\ldots$ with $\lambda_{k}^{2}\to\infty$, and the set of corresponding eigenfunctions $(e_{j})$ forms an orthonormal basis of $L^{2}(M)$. By standard perturbation theory, if $V\in L^{\infty}(M)$, then the eigenvalues of $-\Delta_{g}+V$ are contained in a union of closed disks,
@@
where $\lambda_{k}^{2}$ are the eigenvalues of $-\Delta_{g}$, $r_{k}:=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, and
@@
Under the assumptions of Theorem 1.1, the operator $-\Delta_{g}+V$ can be defined as an $m$-sectorial operator by standard quadratic form methods; see, e.g., [14], or more specifically, [10, Appendix B]. The proof uses Sobolev embedding, as in the case of real-valued $V$, considered in [2, Section 6].
@@ 1.2. Comparison with the Euclidean case
Moreover, if $n\geq 3$ and $q=n/2$, then there exists a constant $c_{n}$ such that there are no eigenvalues outside $[0,\infty)$ if $\|V\|_{L^{n/2}({\mathbb{R}}^{n})}\leq c_{n}$. The one-dimensional bound is due to Abramov–Aslanyan–Davies [1], and the higher dimensional bounds are due to Frank [9]. It is known [3] that the inequality (8) fails for $q>(n+1)/2$, whereas Laptev and Safronov [16] had conjectured its validity up to $q<n$. The limitation to $q\leq(n+1)/2$ essentially comes from the Knapp example in harmonic analysis. It requires that $V$ decay sufficiently fast at infinity in an average sense. On a compact manifold, such a condition is of course meaningless. However, if we consider the $n$-dimensional flat torus $\mathbb{T}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ and rescale by a factor $L>0$, we obtain
@@
Hence, for the rescaled operator, the decay of the potential is now measured relative to the scale $L$. According to Theorem 1.1, its spectrum is contained in the $L^{-2}$ dilation of the set on the right-hand side of (3). In Section 4, we will show that, as $L\to\infty$, all limit points of this set lie in the right-hand side of (8), provided that $n/2<q\leq(n+1)/2$. More precisely, it follows from the proof of Lemma 4.1 that limit points of the first set in (3) (the union of disks) are contained in $[0,\infty)$, while limit points of the second set of (3) are contained in the second set on the right-hand side of (8).
@@ 1.3. Saturation on the sphere and Zoll manifolds
We turn to the optimality of the first set on the right-hand side of (3) when $M=S^{n}$, the round sphere, or more generally, a Zoll manifold. These are manifolds whose geodesics are periodic with a common minimal period. Without loss of generality after rescaling the metric, we may assume that the common minimal period is $2\pi$. By a theorem of Weinstein [23], there is a constant $\alpha=\alpha_{M}\geq 0$ so that all the nonzero eigenvalues of $-\Delta_{g}$ cluster around the values $(k+\alpha)^{2}$, $k\in{\mathbb{N}}$. More precisely, there is a constant $A=A_{M}$ such that each non-zero eigenvalue is contained in a cluster $[(k+\alpha-A/k)^{2},(k+\alpha+A/k)^{2}]$ for some $k\in{\mathbb{N}}$, and the number of eigenvalues in this cluster, counted with multiplicity, is $d_{k}\approx k^{n-1}$. We label the eigenvalues in the $k$-th cluster by $\lambda_{k,j}^{2}$, $j=1,\ldots,d_{k}$.
@@
i) Let $n\geq 2$ and $q>n/2$. For any $\rho>0$, $\theta\in[0,2\pi)$ and $k\gg 1$, there is a potential $V=V_{k,\theta,\rho}\in L^{q}(S^{n})$, $\|V\|_{L^{q}(S^{n})}\approx\rho$, such that
@@
(ii) Theorem 1.3 shows that, on the sphere or more general Zoll manifolds, the radius of the disks in Theorem 1.1, $r_{k}=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, cannot be decreased (up to the value of the constant). More precisely, (3) cannot hold if $r_{k}$ there is replaced by $\epsilon(\lambda_{)}r_{k}$, where ϵ:++:italic-$\epsilon:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is any decreasing function satisfying $\lim_{\lambda\to\infty}\epsilon(\lambda)=0$.
@@ 1.4. Resolvent estimates
Frank [9] proved the Euclidean bound (8) by combining uniform $L^{p}\to L^{p^{\prime}}$ resolvent bounds, established by Kenig, Ruiz, and Sogge [15] in 1987, with the well-known Birman–Schwinger principle. The latter stipulates that $z$ is an eigenvalue of $-\Delta+V$ if and only if $-1$ is an eigenvalue of the compact operator $|V|^{1/2}(-\Delta-z)^{-1}V^{1/2}$. Here and in the following, $V^{1/2}:=V/|V|^{1/2}$. We will follow the same strategy for proving Theorem 1.1. Compared to the Euclidean case, $L^{p}\to L^{p^{\prime}}$ resolvent bounds for the Laplace–Beltrami operator on compact manifolds have a comparatively short history. To make a direct connection to Theorem 1.1 we shall adopt the following convention throughout this paper,
@@ 2. Resolvent estimates
Next, if $f$ has frequencies $\lambda_{j}\leq\sqrt{2A}$, then by Young’s inequality, $\|f\|_{L^{r}(M)}\lesssim_{A}\|f\|_{L^{s}(M)}$ for all $1\leq s\leq r\leq\infty$ (see e.g. [4, Lemma 3.1]). In particular, if $(e_{j})$ is an orthonormal basis of eigenfunctions of $-\Delta_{g}$ in $L^{2}(M)$ with eigenvalues $\lambda_{j}^{2}$ and $E_{j}$ denotes the orthogonal projection $E_{j}f=\langle e_{j},f\rangle e_{j}$, then
@@
In the last inequality, we also used that #$\#\{j:\lambda_{j}\leq\sqrt{2A}\}\lesssim_{A}1$ since the spectrum is discrete. ∎
@@
If $\mathrm{Re}\,\sqrt{z}\leq 10$, then (22) follows from Lemma 2.3, so it remains to consider the case $\mathrm{Re}\,\sqrt{z}>10$. We recall that we are also assuming $\mathrm{Im}\,\sqrt{z}<1$. We will split the resolvent into two pieces, one involving frequencies close to $\mathrm{Re}\,\sqrt{z}$ and the other involving frequencies at least a unit distance from $\mathrm{Re}\,\sqrt{z}$, i.e. $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|\leq\delta$ for the first piece and $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|>\delta$ for the second piece. Here, $0<\delta\leq 1$ is fixed, and we will take $\delta=1/2$ throughout (any other choice can be reduced to this one by scaling the metric). We will prove
@@
The proof of (33) is reduced to the case $|\mathrm{Im}\,\sqrt{z}|\geq 1$ by the following argument (similar to the proof of [4, (2.33)]: Without loss of generality, assume $\mathrm{Im}\,\sqrt{z}\geq 1$. Choose $z_{1}\in{\mathbb{C}}$ with $\mathrm{Re}\,\sqrt{z_{1}}=\mathrm{Re}\,\sqrt{z}$ and $\mathrm{Im}\,\sqrt{z_{1}}=1$. Then
@@
By the Birman–Schwinger principle, $z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{g})$ is an eigenvalue of $-\Delta_{g}+V$ if and only if $-1$ is an eigenvalue of the (compact) Birman–Schwinger operator $K(z):=|V|^{1/2}(-\Delta_{g}-z)^{-1}V^{1/2}$. In particular, if $z$ is an eigenvalue, then $K(z)$ has norm at least one, which implies that
@@
Therefore, (62) implies that $\|V\|_{L^{\frac{n}{2}}(M)}\geq C_{0}^{-1}$. Put differently, $z$ cannot be an eigenvalue if $|\mathrm{Im}\,\sqrt{z}|\geq 1$ and $\|V\|_{L^{\frac{n}{2}}(M)}<C_{0}^{-1}$. Thus, if we set $c:=C_{0}^{-1}/2$, then all eigenvalues of $-\Delta_{g}+V$, $\|V\|_{L^{\frac{n}{2}}(M)}\leq c$, are contained in the region $|\mathrm{Im}\,\sqrt{z}|<1$. Using (22), we obtain the refinement stated in (5). ∎
@@ 3.1. Optimality on $S^{n}$
and that the eigenspace $H_{k}$ corresponding to the eigenvalue $k(k+n-1)$, the space of spherical harmonics of degree $k$, has dimension
@@
Step 1: By the Birman–Schwinger principle (see the first paragraph in the proof of Theorem 1.1), it suffices to find $\chi,\kappa$ as above such that $-\kappa^{-1}$ is an eigenvalue of the compact operator $\chi(-\Delta_{S^{n}}-z_{k,\theta,\rho})^{-1}\chi$, where
@@
Thus, if $q>n/2$ and $k$ is sufficiently large, or if $q=n/2$ and $\rho$ is sufficiently small, we have
@@
Step 2: Since $\chi P_{k}\chi$ is a nonnegative selfadjoint operator of rank at most $d_{k}$, it has a discrete set of nonzero eigenvalues (counted with multiplicity),
(79) $\|\chi P_{k}\chi\|=a_{0}(k)\geq a_{1}(k)\geq\ldots\geq a_{d_{k}}(k)\geq 0,$
@@
(ii) On $S^{n}$, the construction is even more explicit, at least when $2\leq p^{\prime}\leq 2(n+1)/(n-1)$. Sogge [19] showed that in this case, (85) is saturated by the highest weight spherical harmonic $Q_{k}$ of degree $k$. Then $\chi_{k}=|Q_{k}|^{\frac{p}{2q}}$ is completely explicit. Sogge also showed that the dual of (81) is saturated by the $k$-th zonal spherical harmonic $Z_{k}$ for $2(n+1)/(n-1)\leq p^{\prime}\leq\infty$, but $Z_{k}$ cannot be directly used in (85) since it is not in $L^{p}(S^{n})$ for $p\leq 2$.
@@
(93) $\max_{w\in\gamma}\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm nr}(w,\kappa_{0}))\|<1,$
@@
(95) #$\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap U)=\#(\operatorname{spec}{\mathcal{A}_{\rm r}(\cdot,\kappa_{0})}\cap U).$
@@
(96) #$\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap D(z_{{\rm r},0}(\kappa_{0}),C\rho^{2}k^{4\sigma(q)-1}))\geq 1,$

### v2

@@ Abstract.
We prove eigenvalue bounds for Schrödinger operator $\Delta_{g}+V$ on compact manifolds with complex potentials $V$. The bounds depend only on an $L^{q}$-norm of the potential, and they are shown to be optimal, in a certain sense, on the round sphere and more general Zoll manifolds. These bounds are natural analogues of Frank’s [9] results in the Euclidean case.
@@ 1.1. Spectral inclusion
with a complex-valued potential $V$. The unperturbed operator $\Delta_{g}$ is a non-negative, unbounded self-adjoint operator in $L^{2}(M)$, defined with respect to the Riemannian volume measure ${\rm d}x$. Since $\Delta_{g}$ has compact resolvent, its spectrum consists of nonnegative eigenvalues $0=\lambda_{0}^{2}\leq\lambda_{1}^{2}\leq\ldots$ with $\lambda_{k}^{2}\to\infty$, and the set of corresponding eigenfunctions $(e_{j})$ forms an orthonormal basis of $L^{2}(M)$. By standard perturbation theory, if $V\in L^{\infty}(M)$, then the eigenvalues of $\Delta_{g}+V$ are contained in a union of closed disks,
@@
where $\lambda_{k}^{2}$ are the eigenvalues of $\Delta_{g}$, $r_{k}:=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, and
@@
Under the assumptions of Theorem 1.1, the operator $\Delta_{g}+V$ can be defined as an $m$-sectorial operator by standard quadratic form methods; see, e.g., [14], or more specifically, [10, Appendix B]. The proof uses Sobolev embedding, as in the case of real-valued $V$, considered in [2, Section 6].
@@ 1.2. Comparison with the Euclidean case
Moreover, if $n\geq 3$ and $q=n/2$, then there exists a constant $c_{n}$ such that there are no eigenvalues outside $0,\infty)$ if $\|V\|_{L^{n/2}({\mathbb{R}}^{n})}\leq c_{n}$. The one-dimensional bound is due to Abramov–Aslanyan–Davies [1], and the higher dimensional bounds are due to Frank [9]. It is known [3] that the inequality (8) fails for $q>(n+1)/2$, whereas Laptev and Safronov [16] had conjectured its validity up to $q<n$. The limitation to $q\leq(n+1)/2$ essentially comes from the Knapp example in harmonic analysis. It requires that $V$ decay sufficiently fast at infinity in an average sense. On a compact manifold, such a condition is of course meaningless. However, if we consider the $n$-dimensional flat torus $\mathbb{T}^{n}={\mathbb{R}}^{n}/{\mathbb{Z}}^{n}$ and rescale by a factor $L>0$, we obtain
@@
Hence, for the rescaled operator, the decay of the potential is now measured relative to the scale $L$. According to Theorem 1.1, its spectrum is contained in the $L^{-2}$ dilation of the set on the right-hand side of (3). In Section 4, we will show that, as $L\to\infty$, all limit points of this set lie in the right-hand side of (8), provided that $n/2<q\leq(n+1)/2$. More precisely, it follows from the proof of Lemma 4.1 that limit points of the first set in (3) (the union of disks) are contained in $0,\infty)$, while limit points of the second set of (3) are contained in the second set on the right-hand side of (8).
@@ 1.3. Saturation on the sphere and Zoll manifolds
We turn to the optimality of the first set on the right-hand side of (3) when $M=S^{n}$, the round sphere, or more generally, a Zoll manifold. These are manifolds whose geodesics are periodic with a common minimal period. Without loss of generality after rescaling the metric, we may assume that the common minimal period is $2\pi$. By a theorem of Weinstein [23], there is a constant $\alpha=\alpha_{M}\geq 0$ so that all the nonzero eigenvalues of $\Delta_{g}$ cluster around the values $(k+\alpha)^{2}$, $k\in{\mathbb{N}}$. More precisely, there is a constant $A=A_{M}$ such that each non-zero eigenvalue is contained in a cluster $(k+\alpha-A/k)^{2},(k+\alpha+A/k)^{2}]$ for some $k\in{\mathbb{N}}$, and the number of eigenvalues in this cluster, counted with multiplicity, is $d_{k}\approx k^{n-1}$. We label the eigenvalues in the $k$-th cluster by $\lambda_{k,j}^{2}$, $j=1,\ldots,d_{k}$.
@@
i) Let $n\geq 2$ and $q>n/2$. For any $\rho>0$, $\theta\in[0,2\pi)$ and $k\gg 1+\rho^{1/(1-2\sigma(q))}$, there is a potential $V=V_{k,\theta,\rho}\in L^{q}(S^{n})$, $\|V\|_{L^{q}(S^{n})}\approx\rho$, such that
@@
(ii) Theorem 1.3 shows that, on the sphere or more general Zoll manifolds, the radius of the disks in Theorem 1.1, $r_{k}=\|V\|_{L^{q}(M)}(1+\lambda_{k})^{2\sigma(q)}$, cannot be decreased (up to the value of the constant). More precisely, (3) cannot hold if $r_{k}$ there is replaced by $\epsilon(\lambda)r_{k}$, where $\epsilon:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is any decreasing function satisfying $\lim_{\lambda\to\infty}\epsilon(\lambda)=0$.
@@ 1.4. Resolvent estimates
Frank [9] proved the Euclidean bound (8) by combining uniform $L^{p}\to L^{p^{\prime}}$ resolvent bounds, established by Kenig, Ruiz, and Sogge [15] in 1987, with the well-known Birman–Schwinger principle. The latter stipulates that $z$ is an eigenvalue of $\Delta+V$ if and only if 1-1 is an eigenvalue of the compact operator $|V|^{1/2}(-\Delta-z)^{-1}V^{1/2}$. Here and in the following, $V^{1/2}:=V/|V|^{1/2}$. We will follow the same strategy for proving Theorem 1.1. Compared to the Euclidean case, $L^{p}\to L^{p^{\prime}}$ resolvent bounds for the Laplace–Beltrami operator on compact manifolds have a comparatively short history. To make a direct connection to Theorem 1.1 we shall adopt the following convention throughout this paper,
@@ 2. Resolvent estimates
Next, if $f$ has frequencies $\lambda_{j}\leq\sqrt{2A}$, then by Young’s inequality, $\|f\|_{L^{r}(M)}\lesssim_{A}\|f\|_{L^{s}(M)}$ for all $1\leq s\leq r\leq\infty$ (see e.g. [4, Lemma 3.1]). In particular, if $(e_{j})$ is an orthonormal basis of eigenfunctions of $\Delta_{g}$ in $L^{2}(M)$ with eigenvalues $\lambda_{j}^{2}$ and $E_{j}$ denotes the orthogonal projection $E_{j}f=\langle e_{j},f\rangle e_{j}$, then
@@
In the last inequality, we also used that $\{j:\lambda_{j}\leq\sqrt{2A}\}\lesssim_{A}1$ since the spectrum is discrete. ∎
@@
If $\mathrm{Re}\,\sqrt{z}\leq 10$, then (22) follows from Lemma 2.3, so it remains to consider the case $\mathrm{Re}\,\sqrt{z}>10$. We recall that we are also assuming $|\mathrm{Im}\,\sqrt{z}|<1$. We will split the resolvent into two pieces, one involving frequencies close to $\mathrm{Re}\,\sqrt{z}$ and the other involving frequencies at least a unit distance from $\mathrm{Re}\,\sqrt{z}$, i.e. $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|\leq\delta$ for the first piece and $|\lambda_{j}-\mathrm{Re}\,\sqrt{z}|>\delta$ for the second piece. Here, $0<\delta\leq 1$ is fixed, and we will take $\delta=1/2$ throughout (any other choice can be reduced to this one by scaling the metric). We will prove
@@
The proof of (33) is reduced to the case $|\mathrm{Im}\,\sqrt{z}|\geq 1$ by the following argument (similar to the proof of [4, (2.33)]: Choose $z_{1}\in{\mathbb{C}}$ with $\mathrm{Re}\,\sqrt{z_{1}}=\mathrm{Re}\,\sqrt{z}$ and $|\mathrm{Im}\,\sqrt{z_{1}}|=1$. Then
@@
By the Birman–Schwinger principle, $z\in{\mathbb{C}}\setminus\operatorname{spec}(-\Delta_{g})$ is an eigenvalue of $\Delta_{g}+V$ if and only if 1-1 is an eigenvalue of the (compact) Birman–Schwinger operator $K(z):=|V|^{1/2}(-\Delta_{g}-z)^{-1}V^{1/2}$. In particular, if $z$ is an eigenvalue, then $K(z)$ has norm at least one, which implies that
@@
Therefore, (62) implies that $\|V\|_{L^{\frac{n}{2}}(M)}\geq C_{0}^{-1}$. Put differently, $z$ cannot be an eigenvalue if $|\mathrm{Im}\,\sqrt{z}|\geq 1$ and $\|V\|_{L^{\frac{n}{2}}(M)}<C_{0}^{-1}$. Thus, if we set $c:=C_{0}^{-1}/2$, then all eigenvalues of $\Delta_{g}+V$, $\|V\|_{L^{\frac{n}{2}}(M)}\leq c$, are contained in the region $|\mathrm{Im}\,\sqrt{z}|<1$. Using (22), we obtain the refinement stated in (5). ∎
@@ 3.1. Optimality on $S^{n}$
and that the eigenspace $H_{k}$ corresponding to the eigenvalue k(k+n1)k(k+n-1), the space of spherical harmonics of degree $k$, has dimension
@@
Step 1: By the Birman–Schwinger principle (see the first paragraph in the proof of Theorem 1.1), it suffices to find $\chi,\kappa$ as above such that $\kappa^{-1}$ is an eigenvalue of the compact operator $\chi(-\Delta_{S^{n}}-z_{k,\theta,\rho})^{-1}\chi$, where
@@
Thus, if $q>n/2$ and $k\gg 1+\rho^{1/(1-2\sigma(q))}$, or if $q=n/2$ and $\rho$ is sufficiently small, we have
@@
Step 2: Since $P_{k}$ has rank $d_{k}$ and $\chi$ is bounded, $\chi P_{k}\chi$ has rank at most $d_{k}$; in particular, it is a compact operator. Moreover, for any $f\in L^{2}(S^{n})$, $\langle\chi P_{k}\chi f,f\rangle=\|P_{k}\chi f\|^{2}_{L^{2}(S^{n})}\geq 0$, thus $\chi P_{k}\chi$ is self-adjoint and non-negative, having at most $d_{k}$ nonzero eigenvalues $a_{j}(k)$ (counted with multiplicity),
(79) $\|\chi P_{k}\chi\|=a_{0}(k)\geq a_{1}(k)\geq\ldots\geq a_{d_{k}}(k)\geq 0.$
By the spectral theorem for compact, self-adjoint operators, $\chi P_{k}\chi$ has an orthonormal basis of eigenfunctions $\{f_{k,i}\}_{i=1}^{d_{k}}$. If $\{e_{k,i}\}_{i=1}^{d_{k}}$ is an orthonormal basis of $\mathop{\rm Ran}P_{k}$ (thus, $e_{k,i}$ are eigenfuncitons of $\Delta_{g}$ corresponding to eigenvalue $\lambda_{k}^{2}$), then the functions $f_{k,j}$ lie in the linear span of $\{\chi e_{k,i}\}_{i=1}^{d_{k}}$. Moreover, the eigenvalues $a_{j}(k)$ can be computed by diagonalizing the Gram matrix $G=(G_{ij})$, with $G_{ij}=\langle\chi e_{k,i},\chi e_{k,j}\rangle$, $1\leq i,j\leq d_{k}$.
@@
(ii) On $S^{n}$, the construction is even more explicit, at least when $2\leq p^{\prime}\leq 2(n+1)/(n-1)$. Sogge [19] showed that in this case, (85) is saturated by the highest weight spherical harmonic $Q_{k}$ of degree $k$. Then $\chi_{k}=|Q_{k}|^{\frac{p}{2q}}$ is completely explicit. Sogge also showed that the dual of (81) is saturated by the $k$-th zonal spherical harmonic $Z_{k}$ for $2(n+1)/(n-1)\leq p^{\prime}\leq\infty$, but $Z_{k}$ cannot be directly used in (85) since it does not have the stated $L^{2}$ to $L^{p}$ norm ratio in that range of $p$.
@@
(93) $\max_{w\in\gamma}\|\mathcal{A}_{\rm r}(w,\kappa_{0})^{-1}(\mathcal{A}(w,\kappa_{0})-\mathcal{A}_{\rm r}(w,\kappa_{0}))\|<1,$
@@
(95) $\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap U)=\#(\operatorname{spec}{\mathcal{A}_{\rm r}(\cdot,\kappa_{0})}\cap U).$
@@
(96) $\#(\operatorname{spec}{\mathcal{A}(\cdot,\kappa_{0})}\cap D(z_{{\rm r},0}(\kappa_{0}),C\rho^{2}k^{4\sigma(q)-1}))\geq 1,$